# Reasoning to Edit: Hypothetical Instruction-Based Image Editing with Visual Reasoning

### v1

@@ 1 Introduction
The successful deployment of diffusion models [40, 12] in text-to-image (T2I) generation [33, 36, 37, 10] has significantly accelerated the development of instruction-based image editing (IIE) [31]. IIE focuses on performing precise and localized modifications on a source image in response to human commands, thereby enhancing the controllability and accessibility of visual manipulation [7]. Prior works typically utilize the CLIP-based text encoder inherent in T2I diffusion models for instruction embedding [5, 48, 18, 51, 8, 50], which is insufficient for comprehending complex instructions. To address this limitation, recent methods [42, 17, 30, 45, 23, 52, 41] have proposed substituting it with multimodal large language models (MLLMs) [43, 27], enabling richer cross-modal understanding and better alignment with user intent. However, despite the significant progress achieved by these efforts, the following limitations remain underexplored:
@@
(L1) Overlooking hypothetical instructions. Existing IIE methods are primarily designed to handle simple, explicit, and goal-directed instructions (e.g., “remove the dog”), which typically correspond to straightforward editing operations such as adding, replacing, or deleting objects. However, users often begin without a clear editing objective and instead express ambiguous intent through hypothetical instructions (e.g., “What would happen if the ice cube was left in the sun?”). As illustrated in Fig. 1, current methods struggle to interpret and act on such inputs. Successfully addressing these cases requires models to go beyond surface-level edits and perform deeper reasoning about real-world context, physical changes, and the causal or temporal implications of the instruction. Moreover, currently there is no dataset specifically designed for hypothetical instruction-based editing that offers sufficient scale and scenario diversity [17, 46, 19, 29].
@@
(L2) Insufficient reasoning ability. While integrating MLLMs [43, 27] can enhance instruction comprehension and improve alignment with user commands to some extent, these models still lack dedicated mechanisms for deep reasoning over hypothetical instructions (see Fig. 1 results of SmartEdit). We attribute this limitation to the prevailing paradigm’s reliance on coarse-grained features extracted directly from the input image and instruction, which fail to capture the fine-grained semantic cues necessary for implicit reasoning or fully exploit the MLLM’s embedded world knowledge [44]. For instance (Fig. 1(a)), reasoning over such an instruction requires jointly interpreting both visual and textual cues. Visually, elements like the cube’s sharp edges, surface gloss, and surrounding environment reveal its physical state and context. Textually, features such as conditional phrasing, object references, and temporal expressions together suggest a melting process.
@@
To address these limitations, we propose a unified solution: a large-scale hypothetical instruction-based dataset, Reason50K, and a tailored reasoning-aware framework, ReasonBrain. Reason50K comprises diverse hypothetical instructions spanning four reasoning categories: physical, temporal, causal, and story-based reasoning, totaling 51,039 samples. Each sample consists of a source image, a hypothetical instruction, and a corresponding target image that reflects the intended edit. ReasonBrain consists of an MLLM and a diffusion model, augmented with two specialized modules for visual guidance, reasoning, and semantic enrichment: the Fine-grained Reasoning Cue Extraction (FRCE) module and the Cross-Modal Enhancer (CME). The FRCE module extracts detailed reasoning cues through two branches: the visual reasoning branch captures both local and global visual semantics to model spatial relationships and object-level interactions, while the textual reasoning branch identifies key object references and contextual intent from hypothetical instructions, enriched with relevant visual context. These fine-grained features, combined with multi-scale image tokens and textual embeddings, are input to the MLLM alongside learnable tokens to implicitly generate reasoning-aware visual guidance. Finally, the CME enhances these signals via semantic complementarity across modalities, producing well-aligned, semantically rich guidance for diffusion-based image editing. In sum, our contributions are as follows:
@@ 2 Related Work
Instruction-based Image Editing (IIE). IIE [5] aims to train generative models to manipulate a given image based on user-provided instructions. A milestone in this field is InstructPix2Pix (IP2P) [5], which was the first to incorporate natural language instructions into image editing by fine-tuning a text-to-image (T2I) diffusion model on paired image-instruction datasets. Subsequent works have built upon IP2P by introducing novel curated datasets to enhance real-world editing performance and generate high-quality outputs [48, 18, 51, 26, 47]. Others have focused on improving instruction-output alignment by incorporating advanced techniques such as reward learning [50, 4] and multi-task training [38]. Recently, researchers have begun to integrate multimodal large language models (MLLMs) [43, 27] into existing image editing paradigms to enhance the model’s ability to understand complex instructions. We refer to these approaches as MLLM-enhanced methods [7, 17, 23, 45, 52, 42]. For instance, MGIE [7] leverages MLLMs to generate expressive instructions and provide explicit guidance, thereby enhancing editing performance. SmartEdit [17] further introduces a bidirectional interaction module that facilitates mutual understanding between the MLLM output and the input image. Despite recent advances, most existing efforts still rely on direct and explicit instructions (e.g., “Remove the ice”), limiting their ability to handle hypothetical instructions (e.g., “What would happen if the ice cube melted?”) that require deeper reasoning. While MLLMs offer general world knowledge, current frameworks lack mechanisms to extract and utilize fine-grained reasoning details. Our ReasonBrain builds upon the MLLM-enhanced paradigm by introducing dedicated reasoning-aware modules for generating precise visual guidance, with an emphasis on accurately inferring the implicit intent and real-world context embedded in hypothetical instructions.
@@
Reasoning-aware Datasets for Image Editing. To date, only a few works have explored reasoning-aware datasets for image editing [17, 46, 19]. For instance, ReasonEdit [17] was designed primarily for evaluating the reasoning capabilities of image editing models and contains only a small set of textual samples, making it insufficient for training. Moreover, it focuses on object-level reasoning based on explicit instructions, while overlooking reasoning about editing operations themselves. EditWorld [46] aims to inject physical dynamics simulation capabilities into models across both realistic and virtual scenarios, but does not emphasize instruction-level reasoning. ReasonPix2Pix [19] introduces indirect instructions in a descriptive, goal-oriented style but lacks the depth and diversity needed for more advanced hypothetical reasoning. In contrast to these prior efforts, we curate Reason50K, a large-scale dataset specifically tailored for hypothetical instruction reasoning, enabling models to understand and execute complex edits grounded in physical, temporal, causal, and narrative scenarios.
@@ 3.1 Reason50K for Reasoning Injecting
To support our proposed HI-IE task, we construct a novel large-scale dataset Reason50K specifically curated to inject hypothetical instruction-reasoning ability into image editing models. Reason50K contains 51,039 samples, each consisting of an input image, a corresponding hypothetical instruction, and a target edited image. Unlike existing datasets that primarily feature explicit or goal-oriented prompts (e.g., “Remove the animal in the mirror” or “Make it snowy”), our instructions are implicit, open-ended, and reasoning-driven (e.g., “What would happen if the ice cube were left at room temperature?” or “What would this bouquet look like if it were split into two separate rose bouquets?”). This shift introduces a significantly higher level of abstraction and demands real-world understanding. Reason50K is constructed using an inverse-style strategy, where the source image is generated from the target. Specifically, we leverage GPT [1] to produce hypothetical instructions based on user-provided prompts, and employ a diffusion model to generate multiple candidate source images. A hybrid scoring and evaluation scheme is then applied to select the most appropriate image, forming the final image pairs. Due to space limitations, the full data curation process is detailed in Sec. A. Moreover, Reason50K is organized around four distinct reasoning scenarios: Physical Reasoning, Temporal Reasoning, Causal Reasoning, and Story Reasoning, each illustrated with representative examples and instructions in Fig. 2.
| Datasets | Reasoning Instruction | Automatic Generated | Open Domain | #Edits | #Editing Types | Source Example | Instruction | Target Example |
| ReasonPix2Pix [19] |
[…]
40,212 | – |
A colorful insect has landed |
| EditWorld [46] |
[…]
8,674 | 7 |
Shifted her gaze to the left. |
| ReasonEdit [17] |
[…]
219 | – |
Please remove the empty plate. |
| Reason50K (Ours) |
[…]
51,039 | 4 |
What would happen if the cat stood in front of a mirror? |
@@
Reason50K vs. Existing Datasets. Tab. 1 summarizes the key differences between our dataset and existing ones. In comparison: 1) ReasonPix2Pix [19] enhances instruction reasoning by using LLMs to rewrite explicit editing commands into goal-oriented implicit instructions. However, the dataset primarily consists of descriptive instructions and lacks more complex, abstract hypothetical instructions that require deeper contextual and causal reasoning. In addition, it lacks systematic categorization. 2) EditWorld [46] focuses on simulating world dynamics across both real and virtual scenarios. Although it includes some hypothetical instructions, its primary objective is to inject physical and temporal simulation capabilities into image editing models, rather than equipping them with implicit reasoning skills for understanding hypothetical instructions. Additionally, the overall scale of the EditWorld dataset is significantly smaller than ours. 3) ReasonEdit [17] is a small-scale dataset designed primarily for evaluation. It focuses on object-level reasoning from explicit instructions and lacks both diversity and editing operation reasoning. Most importantly, it is not sufficient for model training. In contrast, our Reason50K is the first to provide systematic, large-scale support for training and evaluating hypothetical instruction reasoning across diverse scenarios. Each instruction requires deep reasoning grounded in contextual understanding and world knowledge—spanning physical, causal, temporal, and story-based scenarios—to guide image editing. This goes beyond surface-level transformations or explicit object manipulation. By incorporating carefully crafted hypothetical instructions, the dataset significantly promotes deeper semantic reasoning within image editing models.
@@ 3.2 ReasonBrain
(1) Visual Reasoning Cues Branch (VRCB). This branch aims to extract fine-grained visual cues from both local and global perspectives. Specifically, the local perspective focuses on capturing object parts, textures, and spatial patterns, which are essential for implicit reasoning tasks involving fine object distinctions and subtle appearance changes. Inspired by the MAE framework [9], we first divide the visual features into patches using a patch adapter , and then apply a patch-level feature extractor to obtain localized visual representations. This process is formalized as: . In contrast, the global perspective captures inter-object relationships and contextual information across the entire scene, which are beneficial for reasoning tasks that require an understanding of object interactions, event causality, and broader scene dynamics. We first employ an instance segmentation model (e.g., SAM [20, 34]) to segment objects from the background in the image. Subsequently, a region-level feature extractor is trained to learn holistic semantic representations for each segmented instance. The entire process is formalized as:
@@
(2) Textual Reasoning Cues Branch (TRCB). This branch aims to extract the key object referenced in , serving as a bridge between linguistic intent and visual reasoning. Specifically, we first employ GPT [1] to extract the referenced object from the instruction and serialize it into a structured object token . We then introduce an ID Controller to enhance the model’s ability to perform object-grounded reasoning by facilitating interaction between the object token and the visual reasoning features . This module not only enriches the object token with visual context, enabling the model to reason about the object beyond its textual description, but also aligns the linguistic reference with its corresponding visual entity, helping to resolve ambiguities and prevent semantic drift during generation. The architecture of the ID Controller is illustrated in Fig. 4(a) and is implemented using a cross-attention layer [24] followed by a feed-forward network, formally defined as:
@@
Subsequently, following prior works [17, 52], we concatenate additional learnable tokens with the extracted features and feed them into the MLLM for guidance generation. This operation transforms the implicit reasoning process into a token prediction task, where the MLLM learns to generate the embeddings of these tokens, which serve as visual editing guidance. The process is formalized as:
@@
Here, denotes concatenation, and is a trainable image adapter that maps visual features into the MLLM’s latent space. The output represents the hidden embeddings of the learnable tokens. In addition, we employ a QFormer [22] to align the feature space between the MLLM and the diffusion model, defined as .
@@
Cross-Modal Enhancer. To compensate for the potential loss of visual and textual details in , we introduce a Cross-Modal Enhancer (CME). The CME consists of a visual-oriented enhancer and a textual-oriented enhancer, both implemented using the same bidirectional interaction mechanism. Specifically, each enhancer comprises five hybrid cross-attention blocks, followed by a linear projection layer and a normalization layer (see Fig. 4(b)). These modules reuse intermediate features generated both before and after the MLLM layer, enabling fine-grained semantic enhancement within each modality. Here, we illustrate the process using the visual-oriented enhancer. First, a set of learnable tokens 111The newly introduced symbols are used only in this section to illustrate the visual-oriented enhancer. interacts with (as key and value) via a cross-attention block to produce . In parallel, (query) attends to in a second cross-attention block, yielding . Next, (query) and (key and value) are fused via a third block, followed by a residual connection and normalization to obtain . This refined representation then serves as key and value in a fourth block, interacting with (query). The output is further refined through a final cross-attention with , producing the enhanced visual representation . Similarly, the textual-oriented enhancer is implemented by replacing and with and , respectively. Finally, the CME module outputs four enhanced features, which are passed to the diffusion model together with and a noisy latent for final image generation. Due to space limitations, the training objectives of ReasonBrain are provided in Sec. B.

### v2

@@ 1 Introduction
The successful deployment of diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) in text-to-image (T2I) generation (Ramesh et al., 2022; Rombach et al., 2022b; Saharia et al., 2022; He et al., 2024) has significantly accelerated the development of instruction-based image editing (IIE) (Nguyen et al., 2024a). IIE focuses on performing precise and localized modifications on a source image in response to human commands, thereby enhancing the controllability and accessibility of visual manipulation (Fu et al., 2024). Prior works typically utilize the CLIP-based text encoder inherent in T2I diffusion models for instruction embedding (Brooks et al., 2023; Zhang et al., 2023; Hui et al., 2024; Zhao et al., 2024; Geng et al., 2024; Zhang et al., 2024), which is insufficient for comprehending complex instructions. To address this limitation, recent methods (Tian et al., 2025; Huang et al., 2024; Nguyen et al., 2024b; Wang et al., 2024b; Li et al., 2024; Zhou et al., 2025; Sun et al., 2025) have proposed substituting it with multimodal large language models (MLLMs) (Touvron et al., 2023; Liu et al., 2024b), enabling richer cross-modal understanding and better alignment with user intent. However, despite the significant progress achieved by these efforts, the following limitations remain underexplored:
@@
(L1) Overlooking hypothetical instructions. Existing IIE methods are primarily designed to handle simple, explicit, and goal-directed instructions (e.g., “remove the dog”), which typically correspond to straightforward editing operations such as adding, replacing, or deleting objects. However, users often begin without a clear editing objective and instead express ambiguous intent through hypothetical instructions (e.g., “What would happen if the ice cube was left in the sun?”). As illustrated in Fig. 1, current methods struggle to interpret and act on such inputs. Successfully addressing these cases requires models to go beyond surface-level edits and perform deeper reasoning about real-world context, physical changes, and the causal or temporal implications of the instruction. Moreover, currently there is no dataset specifically designed for hypothetical instruction-based editing that offers sufficient scale and scenario diversity (Huang et al., 2024; Yang et al., 2024; Jin et al., 2024; Meng et al., 2024).
@@
(L2) Insufficient reasoning ability. While integrating MLLMs (Touvron et al., 2023; Liu et al., 2024b) can enhance instruction comprehension and improve alignment with user commands to some extent, these models still lack dedicated mechanisms for deep reasoning over hypothetical instructions (see Fig. 1 results of SmartEdit). We attribute this limitation to the prevailing paradigm’s reliance on coarse-grained features extracted directly from the input image and instruction, which fail to capture the fine-grained semantic cues necessary for implicit reasoning or fully exploit the MLLM’s embedded world knowledge (Wang et al., 2024a). For instance (Fig. 1(a)), reasoning over such an instruction requires jointly interpreting both visual and textual cues. Visually, elements like the cube’s sharp edges, surface gloss, and surrounding environment reveal its physical state and context. Textually, features such as conditional phrasing, object references, and temporal expressions together suggest a melting process.
@@
To address these limitations, we propose a unified solution: a large-scale hypothetical instruction-based dataset, Reason50K, and a tailored reasoning-aware framework, ReasonBrain. Reason50K comprises diverse hypothetical instructions spanning four reasoning categories: physical, temporal, causal, and story-based reasoning, totaling 51,039 samples. Each sample consists of a source image, a hypothetical instruction, and a corresponding target image that reflects the intended edit. ReasonBrain is a hybrid framework that jointly and interactively performs reasoning and editing, thereby overcoming the need for multi-round refinement of textual instructions (Fu et al., 2024). This unified design mitigates reasoning uncertainty and reduces inference time. Specifically, ReasonBrain consists of an MLLM and a diffusion model, augmented with two specialized modules for visual guidance, reasoning, and semantic enrichment: the Fine-grained Reasoning Cue Extraction (FRCE) module and the Cross-Modal Enhancer (CME). The FRCE module extracts detailed reasoning cues through two branches: the visual reasoning branch captures both local and global visual semantics to model spatial relationships and object-level interactions, while the textual reasoning branch identifies key object references and contextual intent from hypothetical instructions, enriched with relevant visual context. These fine-grained features, combined with multi-scale image tokens and textual embeddings, are input to the MLLM alongside learnable tokens to implicitly generate reasoning-aware visual guidance. Finally, the CME enhances these signals via semantic complementarity across modalities, producing well-aligned, semantically rich guidance for diffusion-based image editing. In sum, our contributions are as follows:
@@ 2 Related Work
Instruction-based Image Editing (IIE). IIE (Brooks et al., 2023) aims to train generative models to manipulate a given image based on user-provided instructions. A milestone in this field is InstructPix2Pix (IP2P) (Brooks et al., 2023), which is the first to incorporate natural language instructions into image editing by fine-tuning a text-to-image (T2I) diffusion model on paired image-instruction datasets. Subsequent works have built upon IP2P by introducing novel curated datasets to enhance real-world editing performance and generate high-quality outputs (Zhang et al., 2023; Hui et al., 2024; Zhao et al., 2024; Liu et al., 2024a; Yu et al., 2024). Others have focused on improving instruction-output alignment by incorporating advanced techniques such as reward learning (Zhang et al., 2024; Bai et al., 2024) and multi-task training (Sheynin et al., 2024). Recently, researchers have integrated multimodal large language models (MLLMs) (Touvron et al., 2023; Liu et al., 2024b) into existing image editing paradigms to enhance the model’s ability to understand complex instructions. We refer to these approaches as MLLM-enhanced methods (Fu et al., 2024; Huang et al., 2024; Li et al., 2024; Wang et al., 2024b; Zhou et al., 2025; Tian et al., 2025). For instance, MGIE (Fu et al., 2024) leverages MLLMs to generate expressive instructions and provide explicit guidance, thereby enhancing editing performance. SmartEdit (Huang et al., 2024) further introduces a bidirectional interaction module that facilitates mutual understanding between the MLLM output and the input image. Despite recent advances, most existing efforts still rely on direct and explicit instructions (e.g., “Remove the ice”), limiting their ability to handle hypothetical instructions (e.g., “What would happen if the ice cube melted?”) that require deeper reasoning. While MLLMs offer general world knowledge, current frameworks lack mechanisms to extract and utilize fine-grained reasoning details. Our ReasonBrain builds upon the MLLM-enhanced paradigm by introducing dedicated reasoning-aware modules for generating precise visual guidance, with an emphasis on accurately inferring the implicit intent and real-world context embedded in hypothetical instructions.
@@
Reasoning-aware Datasets for Image Editing. Only a few works have explored reasoning-aware datasets for image editing (Huang et al., 2024; Yang et al., 2024; Jin et al., 2024). ReasonEdit (Huang et al., 2024) is designed primarily for evaluating the reasoning capabilities of image editing models and contains only a small set of textual samples, making it insufficient for training. Moreover, it focuses on object-level reasoning based on explicit instructions, while overlooking reasoning about editing operations themselves. EditWorld (Yang et al., 2024) aims to inject physical dynamics simulation capabilities into models across both realistic and virtual scenarios, but does not emphasize instruction-level reasoning. ReasonPix2Pix (Jin et al., 2024) introduces indirect instructions in a descriptive, goal-oriented style but lacks the depth and diversity needed for more advanced hypothetical reasoning. In contrast to these prior efforts, we curate Reason50K, a large-scale dataset specifically tailored for hypothetical instruction reasoning, enabling models to understand and execute complex edits grounded in physical, temporal, causal, and narrative scenarios.
@@ 3.1 Reason50K for Reasoning Injecting
To support our proposed HI-IE task, we construct a novel large-scale dataset Reason50K specifically curated to inject hypothetical instruction-reasoning ability into image editing models. Reason50K contains 51,039 samples, each consisting of an input image, a corresponding hypothetical instruction, and a target edited image. Unlike existing datasets that primarily feature explicit or goal-oriented prompts (e.g., “Remove the animal in the mirror” or “Make it snowy”), our instructions are implicit, open-ended, and reasoning-driven (e.g., “What would happen if the ice cube were left at room temperature?” or “What would this bouquet look like if it were split into two separate rose bouquets?”). This shift introduces a significantly higher level of abstraction and demands real-world understanding. Reason50K is constructed using an inverse-style strategy, where the source image is generated from the target. Specifically, we leverage GPT (Achiam et al., 2023) to produce hypothetical instructions based on user-provided prompts, and employ a diffusion model to generate multiple candidate source images. A hybrid scoring and evaluation scheme is then applied to select the most appropriate image, forming the final image pairs. Due to space limitations, the full data curation process is detailed in App. C. Moreover, Reason50K is organized around four distinct reasoning scenarios: Physical Reasoning, Temporal Reasoning, Causal Reasoning, and Story Reasoning, each illustrated with representative examples and instructions in Fig. 2.
| Datasets |
#Edits |
Source Example | Instruction | Target Example | ||||||||||
@@
Reason50K vs. Existing Datasets. Tab. 1 summarizes the key differences between our dataset and existing ones. In comparison: 1) ReasonPix2Pix (Jin et al., 2024) enhances instruction reasoning by using LLMs to rewrite explicit editing commands into goal-oriented implicit instructions. However, the dataset primarily consists of descriptive instructions and lacks more complex, abstract hypothetical instructions that require deeper contextual and causal reasoning. In addition, it lacks systematic categorization. 2) EditWorld (Yang et al., 2024) focuses on simulating world dynamics across both real and virtual scenarios. Although it includes some hypothetical instructions, its primary objective is to inject physical and temporal simulation capabilities into image editing models, rather than equipping them with implicit reasoning skills for understanding hypothetical instructions. Additionally, the overall scale of the EditWorld dataset is significantly smaller than ours. 3) ReasonEdit (Huang et al., 2024) is a small-scale dataset designed primarily for evaluation. It focuses on object-level reasoning from explicit instructions and lacks both diversity and editing operation reasoning. Most importantly, it is not sufficient for model training. In contrast, our Reason50K is the first to provide systematic, large-scale support for training and evaluating hypothetical instruction reasoning across diverse scenarios. Each instruction requires deep reasoning grounded in contextual understanding and world knowledge—spanning physical, causal, temporal, and story-based scenarios—to guide image editing. This goes beyond surface-level transformations or explicit object manipulation. By incorporating carefully crafted hypothetical instructions, the dataset significantly promotes deeper semantic reasoning within image editing models.
It is noted that our dataset Reason50K is constructed from synthetic data, as it is difficult to obtain image pairs from videos that support hypothetical editing with clear reasoning structures. In addition, there is currently no standardized benchmark or evaluation protocol for the frame-by-frame assessment of reasoning-based edits. Nevertheless, the generation process of Reason50K is carefully designed to ensure both semantic consistency and high visual quality, thereby enabling the synthetic dataset to effectively support generalizable research.
@@ 3.2 ReasonBrain
(1) Visual Reasoning Cues Branch (VRCB). This branch aims to extract fine-grained visual cues from both local and global perspectives. Specifically, the local perspective focuses on capturing object parts, textures, and spatial patterns, which are essential for implicit reasoning tasks involving fine object distinctions and subtle appearance changes. Inspired by the MAE framework (He et al., 2022), we first divide the visual features into patches using a patch adapter , and then apply a patch-level feature extractor to obtain localized visual representations. This process is formalized as: . In contrast, the global perspective captures inter-object relationships and contextual information across the entire scene, which are beneficial for reasoning tasks that require an understanding of object interactions, event causality, and broader scene dynamics. We first employ an instance segmentation model (e.g., SAM (Kirillov et al., 2023; Ravi et al., 2024)) to segment objects from the background in the image. Subsequently, a region-level feature extractor is trained to learn holistic semantic representations for each segmented instance. The entire process is formalized as:
@@
(2) Textual Reasoning Cues Branch (TRCB). This branch aims to extract the key object referenced in , serving as a bridge between linguistic intent and visual reasoning. Specifically, we first employ GPT (Achiam et al., 2023) to extract the referenced object from the instruction and serialize it into a structured object token . We then introduce an ID Controller to enhance the model’s ability to perform object-grounded reasoning by facilitating interaction between the object token and the visual reasoning features . This module not only enriches the object token with visual context, enabling the model to reason about the object beyond its textual description, but also aligns the linguistic reference with its corresponding visual entity, helping to resolve ambiguities and prevent semantic drift during generation. The architecture of the ID Controller is illustrated in Fig. 4(a) and is implemented using a cross-attention layer (Lin et al., 2022) followed by a feed-forward network, formally defined as:
@@
Subsequently, following prior works (Huang et al., 2024; Zhou et al., 2025), we concatenate additional learnable tokens with the extracted features and feed them into the MLLM for guidance generation. This operation transforms the implicit reasoning process into a token prediction task, where the MLLM learns to generate the embeddings of these tokens, which serve as visual editing guidance. The process is formalized as:
@@
Here, denotes concatenation, and is a trainable image adapter that maps visual features into the MLLM’s latent space. The output represents the hidden embeddings of the learnable tokens. In addition, we employ a QFormer (Li et al., 2023) to align the feature space between the MLLM and the diffusion model, defined as .
@@
Cross-Modal Enhancer. To compensate for the potential loss of visual and textual details in , we introduce a Cross-Modal Enhancer (CME). The CME consists of a visual-oriented enhancer and a textual-oriented enhancer, both implemented using the same bidirectional interaction mechanism. Specifically, each enhancer comprises five hybrid cross-attention blocks, followed by a linear projection layer and a normalization layer (see Fig. 4(b)). These modules reuse intermediate features generated both before and after the MLLM layer, enabling fine-grained semantic enhancement within each modality. Here, we illustrate the process using the visual-oriented enhancer. First, a set of learnable tokens 111The newly introduced symbols are used only in this section to illustrate the visual-oriented enhancer. interacts with (as key and value) via a cross-attention block to produce . In parallel, (query) attends to in a second cross-attention block, yielding . Next, (query) and (key and value) are fused via a third block, followed by a residual connection and normalization to obtain . This refined representation then serves as key and value in a fourth block, interacting with (query). The output is further refined through a final cross-attention with , producing the enhanced visual representation . Similarly, the textual-oriented enhancer is implemented by replacing and with and , respectively. Finally, the CME module outputs four enhanced features, which are passed to the diffusion model together with and a noisy latent for final image generation. Due to space limitations, the training objectives of ReasonBrain are provided in App. D.